# Sparse Probabilistic Synthesis of Quantum Operations

### v1

@@ Abstract
Successful implementations of quantum technologies require protocols and algorithms that use as few quantum resources as possible. Many applications require a desired quantum operation, such as rotation gates in quantum computing or broadband pulses in NMR or MRI applications, that is not feasible to directly implement or would require longer coherence times than achievable. This work develops an approach that enables—at the cost of a modestly increased measurement repetition rate—the exact implementation of such operations. One proceeds by first building a library of a large number of different approximations to the desired gate operation; by randomly selecting these operations according to a pre-optimised probability distribution, one can on average implement the desired operation with a rigorously controllable approximation error. The approach relies on sophisticated tools from convex optimisation to efficiently find optimal probability distributions. A diverse spectrum of applications are demonstrated as (a) exactly synthesising rotations in fault-tolerant quantum computers using only short T-depth circuits and (b) synthesising broadband and band-selective pulses of superior performance in quantum optimal control with (c) further applications in NMR or MRI. The approach is very general and a broad spectrum of practical applications in quantum technologies are explicitly demonstrated.
@@ I Introduction
Here we develop an approach that allows us to implement an ideal, desired quantum operation $\mathcal{U}_{desired}$ that is a central resource in quantum technologies but would otherwise not be possible to directly implement, e.g., due to T-depth limitations in quantum computing, finite rotation-angle resolution in quantum simulators or due to finite offset effects in quantum optimum control, quantum metrology, NMR or MRI applications. The present approach overcomes these limitations by approximating the ideal operation $\mathcal{U}_{desired}$ as a linear combination of a large number of operations $\mathcal{U}_{l}$ that the hardware can natively implement. We start with an offline pre-processing step whereby we carefully construct a library of gate operations $\mathcal{U}_{l}$ and efficiently find a sparse quasiprobability decomposition $\gamma_{l}$ through the use of established tools in convex optimisation. The online, experimental implementation then proceeds by randomly choosing the operations $\mathcal{U}_{l}$ according to probabilities determined by $\gamma_{l}$. This protocol can on average implement the desired quantum operation $\mathcal{U}_{desired}$ through a rigorously controllable, arbitrarily small approximation error.
@@
The present approach resembles to certain quantum error mitigation techniques which, however, address a distinct problem as they mitigate the effect of hardware imperfections [14, 15]. In fact, we will first assume gate operations to be noise-free – while we discuss below that indeed the approach can naturally be combined with quantum error mitigation protocols. Furthermore, Probabilistic Error Cancellation [16, 17, 18] typically applies Clifford recovery operations which would be many orders of magnitude less efficient in the present context as we demonstrate below. In contrast, the present approach uses a large number of operations $\mathcal{U}_{l}$ that are close to the desired operation – this however, makes our classical pre-processing step rather challenging as it requires manipulating ill-conditioned design matrices and thus standard tools from convex optimisation are impractical. For example, we will consider sequences $\mathcal{U}_{l}$ of Clifford and T gates that approximate the same continuous rotation, and are thus very close to each other $\lVert\mathcal{U}_{l}-\mathcal{U}_{k}\rVert\leq 10^{-3}$ resulting in an ill-conditioned optimisation problem.
@@
We address this technical challenge by first posing the decomposition $\gamma_{l}$ as the solution of an underdetermined linear system of equations. Given underdetermined equations may have infinitely many solutions, our aim is to find the optimal one that has a minimal $\lVert\gamma\rVert_{1}$ norm and thus guarantees a minimal measurement overhead of the approach. In fact, an important result in the theory of linear systems is that the minimal L1-norm solution is almost always the sparsest [19, 20] – finding such sparse solutions has become a very well explored and understood field in mathematics with numerous important applications in engineering, signal processing (compressed sensing) and in machine learning [19, 20]. We employ sophisticated Least Angle Regression (LARS) solvers [19, 21] that can achieve arbitrary precision while allowing us to efficiently prepare not only the exact solution but an entire, continuous family of solutions that make a tradeoff between approximation error and L1 norm (measurement overhead).
@@
We demonstrate a diverse spectrum of practical applications in quantum technologies. First, we demonstrate the exact synthesis of continuous-angle rotations in fault-tolerant quantum computers using only short T-depth sequences such that the measurement overhead achieved is more than 4 orders of magnitude lower than with previous techniques. Second, we demonstrate how the present approach enables exactly or approximately synthesising broadband and band-selective pulses in quantum optimum control – which can ultimately lower error rates due to finite-bandwidth effects in native, physical quantum gates used in today’s quantum computers and quantum simulators. Third, we extend these results to NMR and MRI applications where broadband and band-selective pulses are crucial [22]. Finally, we considerably expand prior results on overcoming finite rotation-angle resolution due to discretisation in the classical control electronics.
@@ II.1 Quasiprobability representations
In the following we refer to a unitary operator $U\in\mathbb{C}^{2^{N}\times 2^{N}}$ in terms of its process matrix $\mathcal{U}$ (which acts on a vectorised density matrix isomorphically with conjugation as $\mathcal{U}\mathrm{vec}[\rho]=\mathrm{vec}[U\rho U^{\dagger}]$). To simplify our presentation, we further assume that $\mathcal{U}$ is the Pauli transfer matrix such that $\mathcal{U}\in\mathbb{R}^{2^{2N}\times 2^{2N}}$ is a real matrix [23]. Due to the exponential dimensionality of the process matrices the approach is limited to only few-qubit operations, however, this is sufficient in most practical applications given quantum circuits are typically built of local gate operations (1 qubit and 2 qubit operations).
@@
with the property $\mathbb{E}(\hat{\mathcal{R}}_{desired})=\mathcal{R}_{desired}$.
Refer to Section B.1 for a proof. Similar sampling schemes have been employed in the context of quantum error mitigation whereby the approach is typically called a quasiprobability representation [17, 16, 25] due to the possibly negative coefficients $\gamma_{l}$; While it is not possible to directly inflict the negative weight in a single experiment, one uses the above scheme to estimate expectation values whereby individual measurement outcomes are multiplied in post-processing with the sign $\mathrm{sign}(\gamma_{l})$ of the relevant circuit variant. Negative signs then cause an increase in variance and thus an increased number of circuit repetitions are required for estimating an observable via the L1 norm $\lVert\gamma\rVert_{1}=\sum_{l}|\gamma_{l}|$.
@@
Refer to Section B.2 for a proof that explicitly constructs the estimator $\hat{o}$. The above measurement cost is exponentially larger than having direct access to $\mathcal{U}_{desired}$ in which case a constant number of shots $N_{s}\leq\epsilon^{-2}$ would suffice. Still, the present approach is highly beneficial in practice as long as the number $\nu$ of gates in the circuit is not significantly larger than $\lVert\gamma_{max}\rVert_{1}^{-1}$ – and indeed the present bounds are quite pessimistic and the actual overhead may be orders of magnitude lower [24, 26].
@@ II.3 Convex optimisation
Such minimal l1-norm solutions are a) almost always the sparsest solutions and b) can be solved efficiently using tools from convex optimisation [19, 20]. More specifically, we consider the more general optimisation that additionally allows us to find approximate solutions as
@@
In the present work we specifically construct design matrices that are ill-conditioned which may result in numerical precision issues. For this reason we use a LARS solver that constructs very accurate solutions via the following steps [19, 21]: One starts by solving Eq. 6 in the limit $\lambda\rightarrow\infty$ via the null vector $\underline{\gamma}=(0,0,\cdots 0)^{T}$. Then the value of $\lambda$ is reduced such that a single vector entry is selected that has the highest potential of reducing the error term $\lVert\textbf{R}\underline{\gamma}-\mathrm{U}_{desired}\rVert$ and a new solution vector is provided with only a single non-zero entry. The approach then adds single entries to the vector gradually reducing the value of $\lambda$ until an exact solution is found (or $\lambda=0$ is approached). The significant advantage of this approach is that we construct the entire path along $\lambda$. Thus we can efficiently obtain a solution where the constraint $\lVert\gamma\rVert_{1}=1$ is exactly satisfied, i.e., no overhead, which is very useful in practice as we demonstrate later.
@@
Figure 1: Overrotation (black and red dots) in terms of spherical angles $\theta$ and $\phi$ of Clifford+T sequences of T-depth at most $36$ relative to the targeted continuous angle rotation gate (green dot). A design matrix $\mathbf{R}$ was constructed using $20$ Clifford+T sequences and a further $15$ Clifford rotations (not shown due to their large overrotations). We exactly synthesise the ideal gate (green dot) with a minimal measurement cost $\lVert\gamma\rVert_{1}-1=10^{-6.7}$ that is more than 4 orders of magnitude lower than achievable via prior techniques [16, 27, 32]. This optimal solution is sparse: black dots are rotation gates that receive zero coefficients $\gamma_{l}=0$ while red dots are gates with non-zero coefficients (their opacity illustrates $|\gamma_{l}|$).
@@ III.1 Application 1: Fault-tolerant gate synthesis
Fault-tolerant quantum computers typically achieve universality through decomposing continuous rotation gates into sequences of (relatively) expensive T-gates and (relatively) cheaper Clifford gates; well-established algorithms are available that efficiently decompose any single-qubit rotation into such sequences [33]. However, in early fault-tolerant quantum computers one will make frugal use of T gates – the limited depth of Clifford+T sequences thus introduces a coherent deviation from the desired quantum operation.
@@
We build on the observation that a large number of Clifford+T sequences can be constructed that contain at most $N_{T}$ T gates and approximate the ideal gate $\mathcal{U}_{desired}$ to a precision at least $\epsilon$. For example, we consider a single-qubit $Z$ rotation $R_{z}(0.234234)$ and explicitly construct a library of $20$ different Clifford+T sequences as $\mathcal{U}_{l}$ that achieve a precision (Hilbert-Schmidt distance) at least $\epsilon\leq 10^{-2.4}$ using no more than $36$ T gates. The design matrix is then obtained by appending a further 15 single-qubit Clifford operations to one of the $\mathcal{U}_{l}$ such that $\mathbf{R}\in\mathbb{R}^{35\times 16}$. Fig. 1 (black and red dots) illustrates the overrotation in spherical angles $\theta$ and $\phi$ with respect to the ideal operation (green) when applying elements of the gate library onto an initial quantum state.
@@
By solving the optimisation problem in Eq. 5 for $\lambda\rightarrow 0$ we obtain an exact solution that allows to exactly synthesise the desired Z rotation gate (green dot) at a minimal measurement overhead $\lVert\gamma\rVert_{1}-1=10^{-6.7}$ – this overhead is more than 4 orders of magnitude lower than using Clifford recovery operations only as in [16, 27, 32] in which case $\lVert\gamma\rVert_{1}-1=10^{-2.3}$. Furthermore, Fig. 1 illustrates that indeed our solution is sparse as most gates in the library receive zero coefficients (black dots, $\gamma_{l}=0$) while red dots correspond to elements of the library that receive non-zero coefficients while their opacity is proportional to the absolute values of $\gamma_{l}$.
@@ III.3 Application 3: broadband and band-selective pulses in NMR and MRI
[…]

### v2

@@ Abstract
Successful implementations of quantum technologies require protocols and algorithms that use as few quantum resources as possible. However, many important quantum operations, such as continuous rotation gates in quantum computing or broadband pulses in NMR or MRI applications, can only be implemented approximately using finite quantum resources. This work develops an approach that enables—at the cost of a modestly increased measurement repetition rate—exact implementations on average. One proceeds by first building a library of a large number of different approximations to the desired gate operation; by randomly selecting these operations according to a pre-optimised probability distribution, one can on average implement the desired operation with a rigorously controllable approximation error. The approach relies on sophisticated tools from convex optimisation to efficiently find optimal probability distributions. A diverse spectrum of applications are demonstrated as (a) exactly synthesising rotations in fault-tolerant quantum computers using only low T-count circuits and (b) synthesising broadband and band-selective pulses of superior performance in quantum optimal control with (c) further applications in NMR or MRI. The approach is very general and a broad spectrum of practical applications in quantum technologies are explicitly demonstrated.
@@ I Introduction
Here we develop an approach that allows us to implement an ideal, desired quantum operation $\mathcal{U}_{desired}$ that is a central resource in quantum technologies but would otherwise not be possible to directly implement, e.g., due to T-count limitations in quantum computing, finite rotation-angle resolution in quantum simulators or due to finite offset effects in quantum optimum control, quantum metrology, NMR or MRI applications. The present approach overcomes these limitations by approximating the ideal operation $\mathcal{U}_{desired}$ as a linear combination of a large number of operations $\mathcal{U}_{l}$ that the hardware can natively implement. We start with an offline pre-processing step whereby we carefully construct a library of gate operations $\mathcal{U}_{l}$ and efficiently find a sparse quasiprobability decomposition $\gamma_{l}$ through the use of established tools in convex optimisation. The online, experimental implementation then proceeds by randomly choosing the operations $\mathcal{U}_{l}$ according to probabilities determined by $\gamma_{l}$. This protocol can on average implement the desired quantum operation $\mathcal{U}_{desired}$ through a rigorously controllable, arbitrarily small approximation error.
@@
The present approach resembles to certain quantum error mitigation techniques which, however, address a distinct problem as they mitigate the effect of hardware imperfections [14, 15]. In fact, we will first assume gate operations are noise-free – while we discuss below that indeed the approach can naturally be combined with quantum error mitigation protocols. Furthermore, Probabilistic Error Cancellation [16, 17, 18] typically applies Clifford recovery operations which would be many orders of magnitude less efficient in the present context as we demonstrate below. In contrast, the present approach uses a large number of operations $\mathcal{U}_{l}$ that are close to the desired operation – this however, makes our classical pre-processing step rather challenging as it requires manipulating ill-conditioned design matrices and thus standard tools from convex optimisation are impractical. For example, we will consider sequences $\mathcal{U}_{l}$ of Clifford and T gates that approximate the same continuous rotation, and are thus very close to each other $\lVert\mathcal{U}_{l}-\mathcal{U}_{k}\rVert\leq 10^{-3}$ resulting in an ill-conditioned optimisation problem.
@@
We address this technical challenge by first posing the decomposition $\gamma_{l}$ as the solution of an underdetermined linear system of equations. Given underdetermined equations may have infinitely many solutions, our aim is to find the optimal one that has a minimal $\lVert\gamma\rVert_{1}$ norm and thus guarantees a minimal measurement overhead of the approach. In fact, an important result in the theory of linear systems is that the minimal L1-norm solution is almost always the sparsest [19, 20] and this sparsity is advantageous for the following key reasons. a) the number of different quantum operations used in the experiment does not scale with the (potentially very large) size of the present classical gate library and b) the solution can be found iteratively using a number of steps that does not scale with the size of the (potentially very large) size of the present classical gate library. Finding such sparse solutions has become a very well explored and understood field in mathematics with numerous important applications in engineering, signal processing (compressed sensing) and in machine learning [19, 20]. We employ sophisticated Least Angle Regression (LARS) solvers [19, 21] that can achieve arbitrary precision while allowing us to efficiently prepare not only the exact solution but an entire, continuous family of solutions that make a tradeoff between approximation error and L1 norm (measurement overhead).
@@
We demonstrate a diverse spectrum of practical applications in quantum technologies. First, we demonstrate the exact synthesis of continuous-angle rotations in fault-tolerant quantum computers using only short T-count sequences such that the measurement overhead achieved is more than 4 orders of magnitude lower than with previous techniques. Second, we demonstrate how the present approach enables exactly or approximately synthesising broadband and band-selective pulses in quantum optimum control – which can ultimately lower error rates due to finite-bandwidth effects in native, physical quantum gates used in today’s quantum computers and quantum simulators. Third, we extend these results to NMR and MRI applications where broadband and band-selective pulses are crucial [22]. Finally, we considerably expand prior results on overcoming finite rotation-angle resolution due to discretisation in the classical control electronics.
@@ II.1 Quasiprobability representations
In the following we refer to a unitary operator $U\in\mathbb{C}^{2^{N}\times 2^{N}}$ in terms of its process matrix $\mathcal{U}$, which acts on a vectorised density matrix isomorphically with conjugation as $\mathcal{U}\mathrm{vec}[\rho]=\mathrm{vec}[U\rho U^{\dagger}]$. To simplify our presentation, we further assume that $\mathcal{U}$ is the Pauli transfer matrix such that $\mathcal{U}\in\mathbb{R}^{2^{2N}\times 2^{2N}}$ is a real matrix [23]. The exponential in $N$ size of the process matrix does not pose a significant limitation in practice given the present approach is only applied to design improved local gate operations, i.e., in most practical applications quantum gates typically only act on a small subset of all qubits. Improved general quantum circuits are then obtained as a composition of the improved local, probabilistic gate operations.
@@
with the property $\mathbb{E}(\hat{\mathcal{U}}_{desired})=\mathcal{U}_{desired}$.
Refer to Section B.1 for a proof. Similar sampling schemes have been employed in the context of quantum error mitigation whereby the approach is typically called a quasiprobability representation [17, 16, 25] due to the possibly negative coefficients $\gamma_{l}$; While it is not possible to directly apply a negative weight in a single experiment, one uses the above scheme to estimate expectation values whereby individual measurement outcomes are multiplied in post-processing with the sign $\mathrm{sign}(\gamma_{l})$ of the relevant circuit variant. Negative signs then cause an increase in variance and thus an increased number of circuit repetitions are required for estimating an observable via the L1 norm $\lVert\gamma\rVert_{1}=\sum_{l}|\gamma_{l}|$.
@@
Refer to Section B.2 for a proof that explicitly constructs the estimator $\hat{o}$. The above measurement cost is exponentially larger than having direct access to $\mathcal{U}_{desired}$ in which case a constant number of shots $N_{s}\leq\epsilon^{-2}$ would suffice. Still, the present approach is highly beneficial in practice as long as the number $\nu$ of gates in the circuit is not significantly larger than $\lVert\gamma_{max}\rVert_{1}^{-1}$ – and indeed the present bounds are quite pessimistic as the actual overhead may be orders of magnitude lower [24, 26].
@@ II.3 Convex optimisation
Such minimal l1-norm solutions are a) almost always the sparsest solutions and b) can be solved efficiently using tools from convex optimisation [19, 20]. More specifically, we consider the general optimisation problem that additionally allows us to find approximate solutions as
@@
Figure 1: Overrotation (black and red dots) in terms of spherical angles $\theta$ and $\phi$ of Clifford+T sequences of T count at most $36$ relative to the targeted continuous angle rotation gate (green dot). A design matrix $\mathbf{R}$ was constructed using $20$ Clifford+T sequences and a further $15$ Clifford rotations (not shown due to their large overrotations). We exactly synthesise the ideal gate (green dot) with a minimal measurement cost $\lVert\gamma\rVert_{1}-1=10^{-6.7}$ that is more than 4 orders of magnitude lower than achievable via prior techniques [16, 27, 32]. This optimal solution is sparse: black dots are rotation gates that receive zero coefficients $\gamma_{l}=0$ while red dots are gates with non-zero coefficients (their opacity illustrates $|\gamma_{l}|$).
In the present work we specifically construct design matrices that are ill-conditioned which may result in numerical precision issues. For this reason we use a Least Angle Regression (LARS) solver that constructs very accurate solutions via the following steps [19, 21]: One starts by solving Eq. 6 in the limit $\lambda\rightarrow\infty$ via the null vector $\underline{\gamma}=(0,0,\cdots 0)^{T}$. Then the value of $\lambda$ is reduced such that a single vector entry is selected that has the highest potential of reducing the error term $\lVert\textbf{R}\underline{\gamma}-\mathrm{U}_{desired}\rVert$ and a new solution vector is provided with only a single non-zero entry. The approach then adds single entries to the vector gradually reducing the value of $\lambda$ until an exact solution is found (or $\lambda=0$ is approached). The significant advantage of this approach is that we construct the entire path along $\lambda$. Thus we can efficiently obtain a solution where the constraint $\lVert\gamma\rVert_{1}=1$ is exactly satisfied, i.e., no overhead, which is very useful in practice as we demonstrate later.
@@ III.1 Application 1: Fault-tolerant gate synthesis
Fault-tolerant quantum computers typically achieve universality through decomposing continuous rotation gates into sequences of (relatively) expensive T-gates and (relatively) cheaper Clifford gates; well-established algorithms are available that efficiently decompose any single-qubit rotation into such sequences [33]. However, in early fault-tolerant quantum computers one will make frugal use of T gates – the limited T count of Clifford+T sequences thus introduces a coherent deviation from the desired quantum operation.
@@
We build on the observation that a large number of Clifford+T sequences can be constructed that contain at most $N_{T}$ T gates and approximate the ideal gate $\mathcal{U}_{desired}$ to a precision at least $\epsilon$. For example, we consider a single-qubit $Z$ rotation $R_{z}(0.234234)$ and explicitly construct a library of $20$ different Clifford+T sequences as $\mathcal{U}_{l}$ that achieve a precision (Hilbert-Schmidt distance) at best $10^{-3.4}$ using at most $36$ T gates. The design matrix is then obtained by appending a further 15 single-qubit Clifford operations to one of the $\mathcal{U}_{l}$ such that $\mathbf{R}\in\mathbb{R}^{35\times 16}$. Fig. 1 (black and red dots) illustrates the overrotation in spherical angles $\theta$ and $\phi$ with respect to the ideal operation (green) when applying elements of the gate library onto an initial quantum state.
@@
By solving the optimisation problem in Eq. 5 for $\lambda\rightarrow 0$ we obtain an exact solution that allows to exactly synthesise the desired $R_{z}$ rotation gate (green dot) at a minimal measurement overhead $\lVert\gamma\rVert_{1}-1=10^{-6.7}$ – this overhead is more than 4 orders of magnitude lower than using Clifford recovery operations only as in [16, 27, 32] in which case $\lVert\gamma\rVert_{1}-1=10^{-2.3}$. Furthermore, Fig. 1 illustrates that indeed our solution is sparse as most gates in the library receive zero coefficients (black dots, $\gamma_{l}=0$) while red dots correspond to elements of the library that receive non-zero coefficients while their opacity is proportional to the absolute values of $\gamma_{l}$.
We note that in the present example we used the simple and optimal approach of ref [33] that requires no ancilla qubits and requires $\approx 3.02\log_{2}(1/\epsilon)+1.77$ T gates to achieve precision $\epsilon$. One can reduce the number of T gates to $\approx 1.03\log_{2}(1/\epsilon)+5.75$ via the probabilistic approach of [34] that requires ancilla qubits, resulting in $17$ T gates used in the present example rather than $36$.
Figure 2: Estimated relative total runtime, as a ratio of relative total T counts, with respect to direct synthesis for an increasing number $\nu$ of rotation gates. Even when we take into account its measurement overhead, the present approach requires in total fewer T gates than direct synthesis as long as the number $\nu$ of gates is less than $(\lVert\gamma\rVert_{1}-1)^{-1}$. Furthermore, the present approach has the significant advantage that it guarantees zero approximation error, whereas direct synthesis is approximate and we considered a very permissive circuit-level error $\epsilon_{circ}=10^{-2}$.
Let us compare the end-to-end time complexity of the present approach with direct synthesis assuming the technique of ref. [34]. Assuming $\nu$ continuous rotation gates yields a multiplicative increase of $\exp{(4.0\times 10^{-7}\times\nu)}$ in the number of circuit repetitions via Eq. 3. Given our library consists of rotations of at best $\epsilon_{p}=10^{-3.4}$, we need $1.03\log_{2}(1/\epsilon_{p})+5.75\approx 17$ T gates per rotation which results in a T count of $17\nu$ of a single circuit execution. Then, considering the measurement overhead we obtain a total T count of
$17\nu\exp{(4.0\times 10^{-7}\times\nu)}.$ (7)
In contrast, using direct synthesis with a realistic $\epsilon_{d}=\nu^{-1}10^{-2}$ requires $5.751.03\log_{2}(1/\epsilon_{d})+$ 5.75 T gates per rotation resulting in a total T count of
$12.6\nu+1.03\nu\log_{2}(\nu).$ (8)
We plot the relative total T counts as an estimate of the relative runtime of both approaches for an increasing number $\nu$ of gates in Fig. 2. Indeed, as long as the number $\nu$ of continuous rotation gates in the circuit is smaller than $(\lVert\gamma\rVert_{1}-1)^{-1}$, the present approach achieves a lower total T count which can therefore reduce the runtime of quantum algorithms. However, a significant further advantage of the present approach is that it guarantees zero approximation error $\epsilon=0$, in contrast to direct synthesis for which we assume a permissive circuit level error $\epsilon_{circ}=10^{-2}$ in the present example. While the example in Fig. 2 is limited to circuits with no more than $\approx 10^{7}$ rotation gates, we can straightforwardly extend the approach by building a new library of rotations using a smaller $\epsilon_{p}$ (given in the present example we used $\epsilon_{p}=10^{-3.4}$ which resulted in $\lVert\gamma\rVert_{1}-1=10^{-6.7}$).
@@ III.3 Application 3: broadband and band-selective pulses in NMR and MRI
$\hat{S}(t)_{desired}=\lVert\gamma\rVert_{1}\text{sign}(\gamma_{l})S_{l}(t),$ (12)